# Sums of four generalized polygonal numbers of almost prime length

### v1

@@ 1 Introduction
It is obvious that if $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ is too arbitrary, i.e. $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(p,p,p,p).$ Then $n$ that can be represented is restricted. We hence restrict ourselves to the case where $\prod_{j=1}^{4}\alpha_{j}$ is odd and square-free. To understand the above sum and the relationship with shifted lattice theory, by completing the square, (1.1) is equivalent to
@@ 2.3 Theta series for cosets
Using [1, Proposition 2.1], a straightforward calculation shows that the theta series of cosets of rank $k$ are modular forms of weight $\frac{k}{2}.$ Therefore is splits into a sum of Eisenstein series $E_{X}$ and a cusp form $G_{X}.$ Let $a_{E_{X}}(n)$ and $a_{G_{X}}(n)$ denote the $n$-th Fourier coefficient of $E_{X}$ and $G_{X}$ for any integer $n\geq 0,$ respectively. We have
@@ 3 Eisenstein series components
Shimura [7, Theorem 1.5] showed that the $h$-th Fourier coefficient of the Eisenstein series part of $\Theta_{X}$ can be expressed as the product of local densities. We refer readers to [8, 9] for more details.
@@
$a_{E_{X}}(h)=\frac{(2\pi)^{2}h}{(16d_{L^{\textbf{d}}})^{\frac{1}{2}}\Gamma(2)L(2,\psi)}\cdot\prod_{p\mid e_{1}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}\cdot\prod_{{p\mid h\atop{p\nmid e_{1}}}}\gamma_{p}(2).$ (3.2)
@@
In this subsection, we compute the local densities $b_{p}(h,\lambda_{\textbf{d}},0)$ of $X^{\textbf{d}}$ in order to evaluate $a_{E_{X^{\textbf{d}}}}(h)$ using (3.2).
We first compute the case for $p=2.$
@@
$b_{p}(h,\lambda_{\textbf{d}},0)=\begin{cases}2^{2+\text{min}_{j}\{\text{ord}_{2}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in 4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{2}$,}\\ 0,&\text{otherwise.}\end{cases}$
@@
By [1 Lemma 3.3]. ∎
For odd $p\mid(m-2)$
@@
By [1 Lemma 3.4]. ∎
@@
$I_{p}(n;\phi)=1+(1-\frac{1}{p})\sum_{1\leq t\leq\text{min}(\textbf{t}_{d},\textbf{t}_{n})\atop l_{p}(t)\text{ even}}\delta_{p}(t)p^{\tau_{p}(t)}+\delta_{p}(\textbf{t}_{n}+1)w_{p}p^{\tau_{p}(\textbf{t}_{n}+1)},$
@@
Let $p$ be an odd prime such that $p\nmid(m-2)(m-4),p\mid\prod_{j=1}^{4}d_{j},$ and $p\nmid\prod_{j=1}^{4}\alpha_{j}$
@@
Lower bound for the $\textbf{d}=1$ case. We are now ready to give a lower bound for $a_{E_{X^{\textbf{1}}}(h)}.$
@@
$\begin{split}\beta_{X^{\textbf{p}^{c}},p}(h):=\frac{1}{p^{c_{1}+c_{2}+c_{3}+c_{4}}}&(\frac{b_{p}(h,\lambda_{\textbf{p}^{c}},0)}{b_{p}(h,\lambda_{\textbf{1}},0)})^{\delta_{p\mid(m-2)}}\\ &\cdot(b_{p}(h,\lambda_{\textbf{p}^{c}},0)\cdot\frac{1}{1-\psi(p)p^{-2}}\cdot\gamma_{p}^{-1}(2))^{\delta_{p\nmid 2(m-2),c\neq 0}}\end{split}$ (3.4)
Therefore we have
@@
$\leq\frac{2}{p}$
and similarly for other cases. ∎
@@
$\max\{\beta_{X^{(p,p,p,1)},p}(h),\beta_{X^{(p,p,1,p)},p}(h),\beta_{X^{(p,1,p,p)},p}(h),\beta_{X^{(1,p,p,p)},p}(h)\}\leq\frac{2}{(p-1)^{2}(p+1)}\leq\frac{64}{p^{3}}$
@@
By Theorem 3.4, we have
@@
and similarly for other cases. ∎
@@ 4 Error from Cusp form
Using Lemma 4.1, we have the following bound.

### v2

@@ 1 Introduction
It is obvious that if $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ is too arbitrary, i.e. $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(p,p,p,p),$ then $n$ that can be represented is restricted. We hence restrict ourselves to the case where $\prod_{j=1}^{4}\alpha_{j}$ is odd and square-free. To understand the above sum and the relationship with shifted lattice theory, we complete the square so that (1.1) becomes
@@
Remark.
We further assume that $m-4\not\equiv 0\pmod{3}$ and $m-4\not\equiv 0\pmod{5}$ due to the following technical issue. Suppose we have $m-4\equiv 0\pmod{3}$ or $m-4\equiv 0\pmod{5}$, then $X=2(m-2)x+4-m\not\equiv 0\pmod{p}$ is equivalent to $x\not\equiv 0\pmod{p}$. Therefore the integers n that can be represented as (1.1) with $x_{1},x_{2},x_{3},x_{4}\not\equiv 0\pmod{p}$ is restricted: when $p=3,$ only those $n$ with $n\equiv 0\pmod{p}$, when $p=5$ only those $n$ with $n\not\equiv 0\pmod{p}$.
@@ 2.3 Theta series for cosets
Using [1, Proposition 2.1], a straightforward calculation shows that the theta series of cosets of rank $k$ are modular forms of weight $\frac{k}{2}.$ Therefore it splits into a sum of Eisenstein series $E_{X}$ and a cusp form $G_{X}.$ Let $a_{E_{X}}(n)$ and $a_{G_{X}}(n)$ denote the $n$-th Fourier coefficient of $E_{X}$ and $G_{X}$ for any integer $n\geq 0,$ respectively. We have
@@ 3 Eisenstein series components
Shimura [7, Theorem 1.5] showed that the $h$-th Fourier coefficient of the Eisenstein series part of $\Theta_{X}$ can be expressed as the product of local densities. We refer the readers to [8] and [9] for more details.
@@
$a_{E_{X^{\textbf{d}}}}(h)=\frac{(2\pi)^{2}h}{(16d_{L^{\textbf{d}}})^{\frac{1}{2}}\Gamma(2)L(2,\psi)}\cdot\prod_{p\mid e_{1}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}\cdot\prod_{{p\mid h\atop{p\nmid e_{1}}}}\gamma_{p}(2).$ (3.2)
Furthermore, $t_{p}$ defined in [7, Section 1.6] is equal to 2 in our case. Therefore for any odd prime $p,$ we have
$\gamma_{p}(2)=\begin{cases}\frac{1-(-p^{1-s})^{v+1}}{1+p^{1-s}},&\text{if $\psi(p)\neq 0$ and $c_{p}\in\mathbb{Z}_{p}^{\times}$,}\\ \frac{1-(-p^{1-s})^{v+2}+p^{2-s}(1-(-p^{1-s})^{v})}{(1+p^{-s})(1+p^{1-s})},&\text{if $\psi(p)\neq 0$ and $c_{p}\not\in\mathbb{Z}_{p}^{\times}$},\\ 1+\xi_{p}(c_{p}h)p^{(v+\sigma)(1-s)},&\text{if $\psi(p)=0$}.\end{cases}$
Where the numbers $v,c_{p},\sigma$ and the function $\xi_{p}$ are defined in [7, Section 1.6].
@@
In this subsection, we compute the local densities $b_{p}(h,\lambda_{\textbf{d}},0)$ of $X^{\textbf{d}}$ in order to evaluate $a_{E_{X^{\textbf{d}}}}(h)$ using (3.2). We first require a formula for
$\tau_{\alpha_{j},p,s}(\sigma):=\int_{\mathbb{Z_{p}}}e_{p}(4(m-2)s\sigma\alpha_{j}((m-2)sx^{2}-(m-4)x))\,dx$
for $\sigma\in\mathbb{Q}_{p}.$ Then by [1, Lemma 3.2], we have
Lemma 3.2.
If $\alpha_{j}s\in\mathbb{Z}_{p}$ satisfies $p\mid 2(m-2)\alpha_{j}s,$ then we have
$\tau_{\alpha_{j},p,s}(\sigma)=\begin{cases}1,&\text{if $4(m-2)\alpha_{j}s\sigma\in\mathbb{Z}_{p}$,}\\ 0,&\text{otherwise.}\end{cases}$
We first compute the local density for the case $p=2.$
@@
$b_{2}(h,\lambda_{\textbf{d}},0)=\begin{cases}2^{2+\text{min}_{j}\{\text{ord}_{2}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in 4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{2}$,}\\ 0,&\text{otherwise.}\end{cases}$
@@
Note that we have
$Q(u+\nu)=\sum_{j=1}^{4}4\alpha_{j}(m-2)^{2}(d_{j}x_{j}-\frac{m-4}{2(m-2)})^{2},$ (3.3)
where $\nu=x_{1}(d_{1}e_{1})+x_{2}(d_{2}e_{2})+x_{3}(d_{3}e_{3})+x_{4}(d_{4}e_{4}).$ Then by (3.1) and Lemma 3.2, we have
$b_{2}(h,\lambda_{\textbf{d}},0)$ $=\int_{\mathbb{Q}_{p}}e_{p}(\sigma(\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}-h))\prod_{j=1}^{4}\tau_{\alpha_{j},2,d_{j}}(\sigma)\,d\sigma$
$=\int_{\frac{1}{4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})}\mathbb{Z}_{2}}e_{p}(-8n(m-2)\sigma)\,d\sigma$
$=\begin{cases}2^{2+\text{min}_{j}\{\text{ord}_{2}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in 4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{2}$,}\\ 0,&\text{otherwise.}\end{cases}$
We now compute the local density for odd $p\mid(m-2).$
@@
By (3.1), Lemma 3.2 and (3.3), we have
$b_{p}(h,\lambda_{\textbf{d}},0)$ $=\int_{\mathbb{Q}_{p}}e_{p}(\sigma(\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}-h))\prod_{j=1}^{4}\tau_{\alpha_{j},2,d_{j}}(\sigma)\,d\sigma$
$=\int_{\frac{1}{(m-2)\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})}\mathbb{Z}_{p}}e_{p}(-8n(m-2)\sigma)\,d\sigma$
$=\begin{cases}p^{\text{ord}_{p}(m-2)+\text{min}_{j}\{\text{ord}_{p}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{p}$,}\\ 0,&\text{otherwise.}\end{cases}$
@@
$b_{p}(h,\lambda_{\textbf{d}},0)=1+(1-\frac{1}{p})\sum_{1\leq t\leq\text{min}(\textbf{t}_{d},\textbf{t}_{n})\atop l_{p}(t)\text{ even}}\delta_{p}(t)p^{\tau_{p}(t)}+\delta_{p}(\textbf{t}_{n}+1)w_{p}p^{\tau_{p}(\textbf{t}_{n}+1)},$
@@
Let $p$ be an odd prime such that $p\nmid(m-2)(m-4),p\mid\prod_{j=1}^{4}d_{j},$ and $p\nmid\prod_{j=1}^{4}\alpha_{j}.$
@@
Proof.
For $\mid N_{p}\mid=0,$ we have $\textbf{t}_{d}=1$ and $L_{p}(t)=\emptyset,$ thus $\delta_{p}(t)=1$ and $\tau_{p}(t)=t.$ Therefore
$b_{p}(h,\lambda_{\textbf{d}},0)$ $=1+(1-\frac{1}{p})\sum_{1\leq t\leq\text{min}(\textbf{t}_{d},\textbf{t}_{n})}p^{t}+w_{p}\cdot p^{\textbf{t}_{n}+1}$
$=\begin{cases}p&\text{if $N\in p\mathbb{Z}_{p}$,}\\ 0&\text{if $N\not\in p\mathbb{Z}_{p}$.}\end{cases}$
For $\mid N_{p}\mid=1,$ wlog we assume $1\in N_{p}.$ Then $t_{1}=0.$ Now the computation splits into two cases, which is $t_{1}<\text{min}(\textbf{t}_{d},\textbf{t}_{n})$ or $t_{1}=\text{min}(\textbf{t}_{d},\textbf{t}_{n}).$ For the first case, as we assume $0=t_{1}<\text{min}(\textbf{t}_{d},\textbf{t}_{n}),$ thus $1=\textbf{t}_{d}\leq\textbf{t}_{n},$ hence $\text{min}(\textbf{t}_{d},\textbf{t}_{n})=1$ and $w_{p}=0$. Now as $l_{p}(1)=1,$ the sum in the middle is empty and the local density is
$b_{p}(h,\lambda_{\textbf{d}},0)=1$
For $t_{1}=\text{min}(\textbf{t}_{d},\textbf{t}_{n}),$ we have $\textbf{t}_{d}>\textbf{t}_{n}=0.$ Therefore the local density is bounded by
$0\leq b_{p}(h,\lambda_{\textbf{d}},0)=1+(\frac{u_{1}}{p})(\frac{\textbf{u}_{n}}{p})\leq 2.$
The remaining cases can be compute similarly. ∎
We are now ready to give a lower bound for $a_{E_{X^{\textbf{1}}}(h)}.$
@@
Remark.
Note that (3.4) follows from the fact that for general $\textbf{d}\in\mathbb{N}^{4},$ we have $b_{p}(h,\lambda_{\textbf{d}},0)=b_{p}(h,\lambda_{p^{\text{ord}_{p}\textbf{d}}},0).$
@@
$\begin{split}\beta_{X^{\textbf{p}^{\textbf{c}}},p}(h):=\frac{1}{p^{c_{1}+c_{2}+c_{3}+c_{4}}}&(\frac{b_{p}(h,\lambda_{\textbf{p}^{c}},0)}{b_{p}(h,\lambda_{\textbf{1}},0)})^{\delta_{p\mid(m-2)}}\\ &\cdot(b_{p}(h,\lambda_{\textbf{p}^{\textbf{c}}},0)\cdot\frac{1}{1-\psi(p)p^{-2}}\cdot\gamma_{p}^{-1}(2))^{\delta_{p\nmid 2(m-2),\textbf{c}\neq\textbf{0}}}\end{split}$ (3.5)
where $p^{\textbf{c}}=(p^{c_{1}},p^{c_{2}},p^{c_{3}},p^{c_{4}})$ for $\textbf{c}=(c_{1},c_{2},c_{3},c_{4}).$ Therefore we have
@@
$\leq\frac{2}{p}.$
Other cases can be bound similarly. ∎
@@
$\max\{\beta_{X^{(p,p,p,1)},p}(h),\beta_{X^{(p,p,1,p)},p}(h),\beta_{X^{(p,1,p,p)},p}(h)\\ ,\beta_{X^{(1,p,p,p)},p}(h)\}\leq\frac{2}{(p-1)^{2}(p+1)}\leq\frac{64}{p^{3}}$
@@
By Theorem 3.5, the maximum value of $b_{p}(h,\lambda_{p^{\textbf{c}}},0)$ where $\textbf{c}\in\{(0,0,0,1),(0,0,1,0),(0,1,0,0),(1,0,0,0)\}$ is achieved when $\textbf{t}_{d}=1,t_{1}=t_{2}=0,t_{3}=1,$ where $N_{p}=\{1,2,3\},$ and $\textbf{t}_{n}\geq\textbf{t}_{d}.$ Then
$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=1+(1-\frac{1}{p})\delta_{p}(1)\leq 2-\frac{1}{p}.$
Therefore we have
@@
For $\textbf{c}\in\{(0,0,1,1),(0,1,0,1),(1,0,0,1),(0,1,1,0),(1,0,1,0),(1,1,0,0)\},$ the maximum value of $b_{p}(h,\lambda_{p^{\textbf{c}}},0)$ is achieved when $\textbf{t}_{d}=1,t_{1}=0,t_{2}=1,$ where $N_{p}=\{1,2\},$ $t_{3}=t_{4}=1,$ where $D_{p}=\{3,4\},$ and $\textbf{t}_{n}<\textbf{t}_{d}.$ Then
$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=1+(\frac{u_{1}}{p})(\frac{\textbf{u}_{n}}{p})\leq 2.$
Therefore we have
$\max\{\beta_{X^{(p,p,1,1)},p}(h),\beta_{X^{(p,1,p,1)},p}(h),\beta_{X^{(p,1,1,p)},p}(h),\\ \beta_{X^{(1,p,p,1)},p}(h),\beta_{X^{(1,p,1,p)},p}(h),\beta_{X^{(1,1,p,p)},p}(h)\}\\ \leq\frac{1}{p^{2}}\cdot 2\cdot\frac{p^{3}}{(p-1)^{2}(p+1)}\leq\frac{16}{p^{2}}.$
For $\textbf{c}\in\{(0,1,1,1),(1,0,1,1),(1,1,0,1),(1,1,1,0)\},$ the maximum value of $b_{p}(h,\lambda_{p^{\textbf{c}}},0)$ is achieved when $\textbf{t}_{d}=1,t_{1}=0$ where $N_{p}=\{1\},$ and $\textbf{t}_{n}<\textbf{t}_{d}.$ Then
$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=1+(\frac{u_{1}}{p})(\frac{\textbf{u}_{n}}{p})\leq 2.$
Therefore we have
$\max\{\beta_{X^{(p,p,p,1)},p}(h),\beta_{X^{(p,p,1,p)},p}(h),\beta_{X^{(p,1,p,p)},p}(h)\\ ,\beta_{X^{(1,p,p,p)},p}(h)\}\leq\frac{1}{p^{3}}\cdot 2\cdot\frac{p^{3}}{(p-1)^{2}(p+1)}\leq\frac{64}{p^{3}}$
And finally for $\textbf{c}=(1,1,1,1),$ we have
$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=\begin{cases}p,&\text{for $n\in p\mathbb{Z}_{p}$}\\ 0,&\text{for $n\not\in p\mathbb{Z}_{p}$}\end{cases}$
Therefore we have
@@ 4 Error from Cusp form
We next bound the cuspidal contribution.